OpenStudy (anonymous):

Evaluate: [(30 + 6) - 3^2] ÷ 9 • 2

OpenStudy (anonymous):

[(30 + 6) - 3^2] ÷ 9 • 2 STEP BY STEP: (30 + 6)=36 so you have... [36 - 3^2] ÷ 9 • 2 3^2=9 so you have... [36 - 9] ÷ 9 • 2 36-9=25 so you have... [25] ÷ 9 • 2 [25] ÷ 9 =25/9 so you have... (25/9) • 2 and at the final step... (25/9) • 2 = (25• 2/9) =50/9 = (45+5)/9 = 5 5/9.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

You don't see the answer ?

OpenStudy (anonymous):

6.6 6 1.5 34

OpenStudy (anonymous):

those are the answer choices to the question

OpenStudy (anonymous):

I wrote it clearly. Either your options or your question is incorrect. 5 5/9 (or 5.5555 ) is the correct solution and the only one, for [(30 + 6) - 3^2] ÷ 9 • 2 . Check your question again please.